Suppose a stock had an initial price of $87 per share, paid a dividend of $2.15 per share during the year, and had an ending share price of $78.

a. Compute the percentage total return. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b. What was the dividend yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What was the capital gains yield? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a. percentage total return=(End value-Beginning value+Dividend)/Beginning value

=(78-87+2.15)/87

=(7.87%)(Approx)(Negative).

b. dividend yield=Dividend/Beginning value

=2.15/87

=2.47%(Approx).

c.capital gains yield=(End value-Beginning value)/Beginning value

=(78-87)/87

=(10.34%)(Approx)(Negative).